# SequentialAttention++ for Block Sparsification: Differentiable Pruning Meets Combinatorial Optimization

### v1

@@ 1 Introduction
Pruning methods for neural networks LeCun et al. (1989) replace dense weight matrices by sparse approximations, which offer improved generalization and inference efficiency in terms of storage, energy consumption, and other computational resources. In various common formulations, the problem of computing the best sparse approximation to a dense weight matrix is intractable as it generalizes the sparse linear regression problem, which is known to be NP-hard even to approximate Natarajan (1995); Foster et al. (2015); Gupte & Vaikuntanathan (2021); Price et al. (2022). Despite this fact, a wide variety of techniques have proven to be quite successful in practice. This includes magnitude pruning, $\ell_{1}$ regularization, greedy coordinate descent, sampling, among others.
While earlier works have focused on unstructured (i.e., entrywise) sparsity, which has been an active and fruitful area, researchers have rapidly recognized the importance of structured sparsity, which enforces that the sparse approximation respects certain patterns, such as block structure. These structural constraints often lead to further efficiency gains due to improved hardware utilization Anwar et al. (2017); Pool & Yu (2021); Liu et al. (2022). Our work thus focuses on developing new and improved techniques for structured sparsification of weight matrices, and in particular on block sparsification Ma et al. (2023).
@@ 1.1 Importance scoring and combinatorial optimization
Among importance scoring algorithms, early popular choices have included magnitude pruning Thimm & Fiesler (1995); Han et al. (2015), where the magnitude of each trainable parameter serves as a proxy for its importance, as well as methods based on gradients Karnin (1990); Sanh et al. (2020), Hessians LeCun et al. (1989); Hassibi et al. (1993); Singh & Alistarh (2020); Frantar & Alistarh (2023), and other statistics of the weights. Other works have incorporated $\ell_{1}$ regularization Wen et al. (2016); Yang et al. (2019) to encourage sparsity. More recently, a class of techniques broadly termed differentiable pruning inspired by techniques for differentiable neural architecture search Liu et al. (2019) have increased in popularity, where importance scores and/or soft masks are trained together with the network weights in a differentiable manner Xiao et al. (2019); Voita et al. (2019); Kang & Han (2020); Ramakrishnan et al. (2020); Savarese et al. (2020); Zhang et al. (2022). Variations of this idea use the network weights themselves to represent the “importance scores”, and simply use a transformation of the original network weights Schwarz et al. (2021); Vanderschueren & Vleeschouwer (2023); Cho et al. (2023).
As for the combinatorial optimization aspects of pruning, the use of iterative or greedy procedures has long been explored and is known to improve sparsification quality over “one-shot” uses of importance scoring algorithms LeCun et al. (1989); Hassibi et al. (1993); Ström (1997); Frankle & Carbin (2019). The work of Halabi et al. (2022) gives a theoretical justification of this observation via connections to weakly submodular optimization. Combinatorial optimization algorithms beyond greedy approaches, especially local search methods that improve sparsity patterns via local swaps such as iterative hard thresholding (IHT), have long been known in the submodular optimization literature, and have recently been shown to be extremely effective when combined with magnitude pruning Evci et al. (2020); Peste et al. (2021); Kuznedelev et al. (2023b). The work of Peste et al. (2021) also provides strong theoretical guarantees for their approach, ACDC. Similar ideas have also been termed as “neuroregeneration” in work of Liu et al. (2021).
Given these two highly fruitful approaches to the problem of pruning neural networks, it is natural to ask how recent advances in importance scoring algorithms and combinatorial optimization algorithms can work in concert. Our work gives a thorough investigation of this question from both theoretical and empirical perspectives.
@@ 1.2 Theoretical results
On the theoretical side, we present an investigation of differentiable pruning techniques for block sparsification when the objective function $\mathcal{L}:\mathbb{R}^{n}\to\mathbb{R}$ is strictly convex and differentiable. Note that this already captures several interesting problems where block sparsification of weight matrices is desired, such as multinomial logistic regression and multiple response linear regression. We take the $n$ variables of our objective function to be partitioned into disjoint groups $\{T_{i}\}_{i=1}^{t}$ where $T_{i}\subseteq[n]$ and possibly have varying size. For instance, in the context of block sparsification, $\mathcal{L}$ could correspond to the multinomial logistic regression objective function with $K$ classes and $d$ features, and the $n=Kd$ variables could be partitioned into $t$ blocks $T_{1},T_{2},\dots,T_{t}$. Furthermore, we will also consider an $\ell_{2}$ regularization term on the parameters $\boldsymbol{\beta}$, that is, we study variants of the problem $\min_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\boldsymbol{\beta})+\lambda\lVert\boldsymbol{\beta}\rVert_{2}^{2}$. Note that explicit $\ell_{2}$ regularization is a standard component of machine learning architectures, and also appears implicitly whenever a loss function is optimized with gradient descent Shalev-Shwartz (2012), with the regularization parameter $\lambda$ being controlled by learning rate parameters and early stopping Suggala et al. (2018).
Our contributions are twofold: (1) we show that a wide variety of differentiable pruning techniques can all be understood as an implementation of nonconvex regularization that generalizes the group LASSO, and (2) we show that a wide class of nonconvex regularizers give a unique sparse global minimum that coincides with the unique sparse global minimum of a corresponding group LASSO problem.
@@ 1.2.1 Differentiable pruning as nonconvex regularization
For the first item, we observe that if we minimize the loss $\mathcal{L}$ with each of the variable groups $\boldsymbol{\beta}|_{T_{i}}$ for $i\in[t]$ replaced by a “masked” version $q(\mathbf{w}_{i})\boldsymbol{\beta}|_{T_{i}}$, and with regularization on $\mathbf{w}$ and $\boldsymbol{\beta}$, then this problem is equivalent to another optimization problem that simply optimizes $\mathcal{L}$ with a different, and often sparsity-inducing, regularizer. A basic version of this observation already appears in works of Hoff (2017); Axiotis & Yasuda (2023), where it is shown that if the masks $q$ are just the identity, then we recover the usual group LASSO problem, that is,
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\frac{\lambda}{2}\left\lparen\lVert\mathbf{w}\rVert_{2}^{2}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen$ )
[…]
@@
where $\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t}$ denotes the concatenation of the “masked” groups $\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}$ for $i\in[t]$. We generalize this observation and show that this framework also applies to other ideas popular in the differentiable pruning literature, such as applying $\ell_{1}$ regularization on the masks $\mathbf{w}$ to induce sparsity Yang et al. (2019) or applying softmax-type masks such as $\exp(\mathbf{w}_{i})$ Yasuda et al. (2023). We note that prior to our work, there was no theoretical understanding on the value of applying such techniques in the context of differentiable pruning.
@@
We also show that similar ideas apply to differentiable pruning techniques that use the network weights themselves as importance scores Schwarz et al. (2021); Cho et al. (2023). Here, the basic observation is that if one optimizes a loss function $\mathcal{L}$ with variables $\boldsymbol{\beta}$ replaced by the (signed) entrywise square $\boldsymbol{\beta}\odot\boldsymbol{\beta}$, then this results in a “rich get richer” dynamic where large weights evolve to be larger while smaller weights are driven down to zero, resulting in sparse solutions. It is known that this idea has connections to exponentiated gradient descent which also results in sparse solutions Vaskevicius et al. (2019); Amid & Warmuth (2020a, b). However, prior work only handles entrywise sparsity and does not address the question of structured pruning. We note that, in fact, these ideas can also be understood in the framework of sparsity-inducing regularizers, even in the group setting. Here, we show that “masking” each of the variable groups $\boldsymbol{\beta}|_{T_{i}}$ by its $\ell_{2}$ norm $\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}$ gives a natural group generalization of this technique, and that this gives an optimization problem that is again equivalent to the group LASSO problem
$\min_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}\cdot\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\lambda\lVert\boldsymbol{\beta}\rVert_{2}^{2}$
[…]
@@ 1.2.2 Unique sparse global minima
For instance, some popular choices of $q$ include the absolute value $q(x)=\lvert x\rvert$, $p$-th powers $q(x)=\lvert x\rvert^{p}$ for $p<1$, or logarithmic regularizers such as $q(x)=\log(1+x)$. In general, the class of such $q$ (strictly) contains the set of all concave functions $q$ that vanish at the origin. Note that the form of (1) slightly differs from the usual form of nonconvex regularizers, as it applies $q^{-1}$ on the sum $\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})$ rather than taking the regularizer to just be $\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})$. This does not substantially change the nature of the optimization problem as it is the Lagrangian dual for the same constraint. Our main result of this section is Theorem 1.1, which relates the group $q$-regularized objective (1) to the following corresponding group LASSO objective:
@@
We make several remarks about Theorem 1.1. First, we justify why the assumption of the theorem is not vacuous: that is, we explain why the group LASSO objective (2) has sparse solutions. Indeed, in recent work of Axiotis & Yasuda (2023) the following are shown if $\mathcal{L}$ is strongly convex and differentiable:
@@
  • If $\lambda<\tau$ is sufficiently close to $\tau$, then (1) has a unique global minimizer with group sparsity $1$.
Thus, when $\lambda$ is sufficiently large, Theorem 1.1 indeed establishes that (1) has a unique sparse global minimum.
Furthermore, Axiotis & Yasuda (2023) also show that the above global minimizer of the group LASSO problem (2) with group sparsity $1$ is supported on a group $T_{i}$ that maximizes $\lVert\nabla\mathcal{L}(0)|_{T_{i}}\rVert_{2}$, that is, it selects the group of variables that locally provides the largest improvement in the objective function cost. Repeatedly alternating between selecting such a feature and re-optimizing over the support is an algorithm known as the group orthogonal matching pursuit (group OMP), and has provable guarantees for group sparse convex optimization when $\mathcal{L}$ satisfies the restricted strong convexity and restricted smoothness properties Axiotis & Yasuda (2023). It is also shown that a related local search algorithm known as group orthogonal matching pursuit with replacement (group OMPR) also applies in this context, which has improved guarantees.
@@ 1.3 Empirical results: SequentialAttention++
We now apply our theoretical insights of combining differentiable pruning and combinatorial optimization to develop a novel algorithm for block neural network pruning, which we call SequentialAttention++. SequentialAttention++ is primarily a fusion of two prior techniques: Sequential Attention, a feature selection technique based on differentiable pruning developed in work of Yasuda et al. (2023), and ACDC, which is a highly effective stochastic adaptation of the classic iterative hard thresholding (IHT) algorithm Peste et al. (2021).
Sequential Attention Yasuda et al. (2023) is an algorithm for feature selection on neural networks, which works by introducing a softmax mask that is trained together with the neural network weights. Each of the $n$ input features is scaled by a differentiable mask $A_{i}=\exp(L_{i})/\sum_{j=1}^{n}\exp(L_{j})$ for a vector $L\in\mathbb{R}^{n}$ of logits. Note that our theoretical results on differentiable pruning, and in particular Lemma 2.1, suggests that this roughly corresponds to performing a log-sum regularization on the corresponding weights for these features. We first extend this to the block sparsification setting by instead scaling each block of weights to prune by a similar softmax mask (see Figure 1). Note that in this new setting, Lemma 2.1 shows that this corresponds to a group log-sum regularization on each of the blocks.
@@
We then use this differentiable pruning technique as part of a local search procedure inspired by ACDC Peste et al. (2021). In the originally proposed ACDC algorithm, the neural network is trained in multiple phases, where the phases alternate between a “dense” training phase and a “sparse” training phase. During the dense phases, the weights are trained in the standard way, whereas in the sparse phases, only a sparse set of weights corresponding to the top $k$ weights at the beginning of the phase (i.e., chosen by magnitude pruning) are used. The idea here is that if a suboptimal sparse support is selected during the sparse phase, then this support can be modified during the dense phase. We note that one of the weaknesses of this algorithm is the use of the weight magnitudes as a proxy for the importance of the weights, whereas improved parameter importance estimation is possible by introducing the use of differentiable pruning techniques. Thus in our SequentialAttention++ algorithm, we modify the ACDC algorithm by training a softmax mask together with the neural network weights during the dense phase as in Figure 1, and then using the softmax mask to select a sparse support during the sparse phases. Our theoretical results establish provable guarantees for a slightly modified version of this algorithm, by showing that log-sum regularization can be integrated with a similar local search algorithm that alternates between dropping small weights from the support, selecting weights via regularization, and optimizing on the new support (see Theorem B.3 and Appendix B).

### v2

@@ 1 Introduction
Pruning methods for neural networks (LeCun et al., 1989) replace dense weight matrices by sparse approximations, which offer improved generalization and inference efficiency in terms of storage, energy consumption, and other computational resources. In various common formulations, the problem of computing the best sparse approximation to a dense weight matrix is intractable as it generalizes the sparse linear regression problem, which is known to be NP-hard even to approximate (Natarajan, 1995; Foster et al., 2015; Gupte and Vaikuntanathan, 2021; Price et al., 2022). Despite this fact, a wide variety of techniques have proven to be quite successful in practice. This includes magnitude pruning, $\ell_{1}$ regularization, greedy coordinate descent, sampling, among others.
While earlier works have focused on unstructured (i.e., entrywise) sparsity, which has been an active and fruitful area, researchers have rapidly recognized the importance of structured sparsity, which enforces that the sparse approximation respects certain patterns, such as block structure. These structural constraints often lead to further efficiency gains due to improved hardware utilization (Anwar et al., 2017; Pool and Yu, 2021; Liu et al., 2022). Our work thus focuses on developing new and improved techniques for structured sparsification of weight matrices, and in particular on block sparsification (Ma et al., 2023), which allow for a balance between performance gains from hardware utilization and reduced computation due to sparsity (Gale et al., 2023).
@@ 1.1 Importance scoring and combinatorial optimization
Among importance scoring algorithms, early popular choices have included magnitude pruning (Thimm and Fiesler, 1995; Han et al., 2015), where the magnitude of each trainable parameter serves as a proxy for its importance, as well as methods based on gradients (Karnin, 1990; Sanh et al., 2020), Hessians (LeCun et al., 1989; Hassibi et al., 1993; Singh and Alistarh, 2020; Frantar and Alistarh, 2023), and other statistics of the weights. Other works have incorporated $\ell_{1}$ regularization (Wen et al., 2016; Yang et al., 2019) to encourage sparsity. More recently, a class of techniques broadly termed differentiable pruning inspired by techniques for differentiable neural architecture search (Liu et al., 2019) have increased in popularity, where importance scores and/or soft masks are trained together with the network weights in a differentiable manner (Xiao et al., 2019; Voita et al., 2019; Kang and Han, 2020; Ramakrishnan et al., 2020; Savarese et al., 2020; Zhang et al., 2022). Variations of this idea use the network weights themselves to represent the “importance scores”, and simply use a transformation of the original network weights (Schwarz et al., 2021; Vanderschueren and Vleeschouwer, 2023; Cho et al., 2023).
As for the combinatorial optimization aspects of pruning, the use of iterative or greedy procedures has long been explored and is known to improve sparsification quality over “one-shot” uses of importance scoring algorithms (LeCun et al., 1989; Hassibi et al., 1993; Ström, 1997; Frankle and Carbin, 2019). The work of Halabi et al. (2022) gives a theoretical justification of this observation via connections to weakly submodular optimization. Combinatorial optimization algorithms beyond greedy approaches, especially local search methods that improve sparsity patterns via local swaps such as iterative hard thresholding (IHT), have long been known in the submodular optimization literature, and have recently been shown to be extremely effective when combined with magnitude pruning (Evci et al., 2020; Peste et al., 2021; Kuznedelev et al., 2023b; Benbaki et al., 2023). The work of Peste et al. (2021) also provides strong theoretical guarantees for their approach, ACDC. Similar ideas have also been termed as “neuroregeneration” in work of Liu et al. (2021).
Given these two highly fruitful approaches to the problem of pruning neural networks, it is natural to ask how recent advances in importance scoring algorithms and combinatorial optimization algorithms can work in concert. We investigate this question from both theoretical and empirical perspectives.
@@ 1.2 Theoretical results
We first present a theoretical investigation of differentiable pruning techniques for block sparsification when the objective function $\mathcal{L}:\mathbb{R}^{n}\to\mathbb{R}$ is strictly convex and differentiable. This already captures several interesting problems where block sparsification of weight matrices is desired, such as multinomial logistic regression and multiple response linear regression. We take the $n$ variables of our objective function to be partitioned into disjoint groups $\{T_{i}\}_{i=1}^{t}$ where $T_{i}\subseteq[n]$ and possibly have varying size. For instance, in the context of block sparsification, $\mathcal{L}$ could correspond to the multinomial logistic regression objective function with $K$ classes and $d$ features, and the $n=Kd$ variables could be partitioned into $t$ blocks $T_{1},T_{2},\dots,T_{t}$. Furthermore, we will also consider an $\ell_{2}$ regularization term on the parameters $\boldsymbol{\beta}$, that is, we study variants of the problem $\min_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\boldsymbol{\beta})+\lambda\lVert\boldsymbol{\beta}\rVert_{2}^{2}$. Note that explicit $\ell_{2}$ regularization is a standard component of machine learning architectures, and also appears implicitly whenever a loss function is optimized with gradient descent (Shalev-Shwartz, 2012), with the regularization parameter $\lambda$ being controlled by learning rate parameters and early stopping (Suggala et al., 2018).
Our contributions are twofold: (1) we show that a wide variety of differentiable pruning techniques can all be understood as an implementation of nonconvex regularization that generalizes the group LASSO, and (2) we show that a wide class of nonconvex regularizers give a unique $1$-sparse global minimum that coincides with the unique 1-sparse global minimum of a corresponding group LASSO problem. These two results together establish that many differentiable pruning techniques work simply by identifying the same $1$-sparse solution as the group LASSO. In turn, it is known that the $1$-sparse solution found by the group LASSO is the variable block with the largest squared gradient (Axiotis and Yasuda, 2023), which is equivalent to the Orthogonal Matching Pursuit (Pati et al., 1993; Shalev-Shwartz et al., 2010; Liberty and Sviridenko, 2017; Elenberg et al., 2018) when applied sequentially (see Appendix B). Thus together, these results make progress towards understanding the inner workings of modern differentiable pruning methods.
@@ 1.2.1 Differentiable pruning as nonconvex regularization
For our first contribution, we observe that if we minimize the loss $\mathcal{L}$ with each of the variable groups $\boldsymbol{\beta}|_{T_{i}}$ for $i\in[t]$ replaced by a “masked” version $q(\mathbf{w}_{i})\boldsymbol{\beta}|_{T_{i}}$, and with regularization on $\mathbf{w}$ and $\boldsymbol{\beta}$, then this problem is equivalent to another optimization problem that simply optimizes $\mathcal{L}$ with a different, and often sparsity-inducing, regularizer. A basic version of this observation already appears in works of Hoff (2017); Axiotis and Yasuda (2023), where it is shown that if the masks $q$ are just the identity, then we recover the usual group LASSO problem, that is,
@@
$\min_{\mathbf{w}\in\mathbb{R}^{t},\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t})+\frac{\lambda}{2}\left\lparen\lVert\mathbf{w}\rVert_{2}^{2}+\lVert\boldsymbol{\beta}\rVert_{2}^{2}\right\rparen=\min_{\boldsymbol{\beta}\in\mathbb{R}^{n}}\mathcal{L}(\boldsymbol{\beta})+\lambda\sum_{i=1}^{t}\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}$
@@
where $\{\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}\}_{i=1}^{t}$ denotes the concatenation of the “masked” groups $\mathbf{w}_{i}\boldsymbol{\beta}|_{T_{i}}$ for $i\in[t]$. We generalize this observation and show that this framework also applies to other ideas popular in the differentiable pruning literature, such as applying $\ell_{1}$ regularization on the masks $\mathbf{w}$ to induce sparsity (Yang et al., 2019) or applying softmax-type masks such as $\exp(\mathbf{w}_{i})$ (Yasuda et al., 2023). We note that prior to our work, there was little theoretical understanding on the value of applying such techniques in the context of differentiable pruning.
@@
We also apply similar ideas to differentiable pruning techniques that use the network weights themselves as importance scores (Schwarz et al., 2021; Cho et al., 2023). Here, the basic observation is that if one optimizes a loss function $\mathcal{L}$ with variables $\boldsymbol{\beta}$ replaced by the (signed) entrywise square $\boldsymbol{\beta}\odot\boldsymbol{\beta}$, then this results in a “rich get richer” dynamic where large weights evolve to be larger while smaller weights are driven to zero, resulting in sparse solutions. This idea also has connections to exponentiated gradient descent which also results in sparse solutions (Vaskevicius et al., 2019; Amid and Warmuth, 2020a, b). However, prior work only handles entrywise sparsity and does not address the question of structured pruning. We show that these ideas can also be understood in the framework of sparsity-inducing regularizers, even in the group setting. Here, we show that “masking” each of the variable groups $\boldsymbol{\beta}|_{T_{i}}$ by its $\ell_{2}$ norm $\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2}$ gives a natural group generalization of this technique, and that this gives an optimization problem that is again equivalent to the group LASSO.
@@ 1.2.2 Unique sparse global minima
For instance, some popular choices of $q$ include the absolute value $q(x)=\lvert x\rvert$, $p$-th powers $q(x)=\lvert x\rvert^{p}$ for $p<1$, or logarithmic regularizers such as $q(x)=\log(1+x)$. In general, the class of such $q$ (strictly) contains the set of all concave functions $q$ that vanish at the origin. Note that the form of (1) slightly differs from the usual form of nonconvex regularizers, as it applies $q^{-1}$ on the sum $\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})$ rather than taking the regularizer to just be $\sum_{i=1}^{t}q(\lVert\boldsymbol{\beta}|_{T_{i}}\rVert_{2})$. This does not substantially change the nature of the optimization problem as it is the Lagrangian dual for the same constraint. The main result of this section is Theorem 1.1, which relates the group $q$-regularized objective (1) to the following corresponding group LASSO objective:
@@
We make several remarks about Theorem 1.1. First, we justify why the assumption of the theorem is not vacuous: that is, we explain why the group LASSO objective (2) has sparse solutions. In recent work of Axiotis and Yasuda (2023) the following are shown if $\mathcal{L}$ is strongly convex and differentiable:
@@
  • If $\lambda<\tau$ is sufficiently close to $\tau$, then (1) has a unique $1$-sparse global minimizer.
Thus, when $\lambda$ is large enough, Theorem 1.1 establishes that (1) has a unique sparse global minimum.
Furthermore, Axiotis and Yasuda (2023) also show that the above global minimizer of the group LASSO problem (2) with group sparsity $1$ is supported on a group $T_{i}$ that maximizes $\lVert\nabla\mathcal{L}(0)|_{T_{i}}\rVert_{2}$, that is, it selects the group of variables that locally provides the largest improvement in the objective function cost. Repeatedly alternating between selecting such a feature and re-optimizing over the support is an algorithm known as the group orthogonal matching pursuit (group OMP), and has provable guarantees for group sparse convex optimization when $\mathcal{L}$ satisfies the restricted strong convexity and restricted smoothness properties (Axiotis and Yasuda, 2023). It is also shown that a related local search algorithm known as group orthogonal matching pursuit with replacement (group OMPR) also applies in this context, which has improved guarantees.
@@ 1.3 Empirical results: SequentialAttention++
We now apply our theoretical insights of combining differentiable pruning and combinatorial optimization to develop a novel algorithm for block neural network pruning, which we call SequentialAttention++. SequentialAttention++ is primarily a fusion of two prior techniques: Sequential Attention, a feature selection technique based on differentiable pruning developed in work of Yasuda et al. (2023), and ACDC, which is a highly effective stochastic adaptation of the classic iterative hard thresholding (IHT) algorithm (Peste et al., 2021) from the combinatorial optimization literature.
Sequential Attention (Yasuda et al., 2023) is an algorithm for feature selection on neural networks, that introduces a softmax mask that is trained together with the neural network weights. Each of the $n$ input features is scaled by a differentiable mask $A_{i}=\exp(L_{i})/\sum_{j=1}^{n}\exp(L_{j})$ for a vector $L\in\mathbb{R}^{n}$ of logits. Note that our theoretical results on differentiable pruning, and in particular Lemma 2.1, suggests that this roughly corresponds to performing a log-sum regularization on the corresponding weights for these features. We first extend this to the block sparsification setting by instead scaling each block of weights to prune by a similar softmax mask (see Figure 1). Note that in this new setting, Lemma 2.1 shows that this corresponds to a group log-sum regularization on each of the blocks.
@@
We then use this differentiable pruning technique as part of a local search procedure inspired by ACDC (Peste et al., 2021). In the originally proposed ACDC algorithm, the neural network is trained in multiple phases, where the phases alternate between a “dense” training phase and a “sparse” training phase. During the dense phases, the weights are trained in the standard way, whereas in the sparse phases, only a sparse set of weights corresponding to the top $k$ weights at the beginning of the phase (i.e., chosen by magnitude pruning) are used. The idea here is that if a suboptimal sparse support is selected during the sparse phase, then this support can be modified during the dense phase. We note that one of the weaknesses of this algorithm is the use of the weight magnitudes as a proxy for the importance of the weights, whereas improved parameter importance estimation is possible by introducing differentiable pruning techniques. Thus in our SequentialAttention++ algorithm, we modify the ACDC algorithm by training a softmax mask together with the neural network weights during the dense phase as in Figure 1, and then using the softmax mask to select a sparse support during the sparse phases. Our theoretical results establish provable guarantees for a slightly modified version of this algorithm, by showing that log-sum regularization can be integrated with a similar local search algorithm that alternates between dropping small weights from the support, selecting weights via regularization, and optimizing on the new support (see Theorem B.3 and Appendix B).